OpenStudy (anonymous):

cosx^2x-1 divided by cosx+1

myininaya (myininaya):

oops i mean cosx^(2x)?

OpenStudy (anonymous):

no, cosine squared x minus one divided by cosine x plus one

OpenStudy (anonymous):

lol

myininaya (myininaya):

(cosx)^2-1?

OpenStudy (anonymous):

the exponent 2 is after the cos lol

myininaya (myininaya):

\[\frac{\cos^2x-1}{cosx+1}=\frac{(cosx+1)(cosx-1)}{cosx+1}\]

OpenStudy (anonymous):

So you have: \[\frac{\cos^2(x)-1}{\cos(x)+1}\]

OpenStudy (anonymous):

so = cosx+1

OpenStudy (anonymous):

ah the method of myininaya!

myininaya (myininaya):

\[(cosx)^2=\cos^2x\] just so you know these are the same

OpenStudy (anonymous):

The top is a difference of squares which factors giving: \[\frac{(\cos(x)-1)(\cos(x)+1)}{cos(x)+1}\rightarrow \cos(x)+1\]

myininaya (myininaya):

the answer is cosx-1

OpenStudy (anonymous):

Yeah, I mistyped, should read cos(x)-1

OpenStudy (anonymous):

motherofgod :( :( :( is it cosx-1 or cosx+1 ?

myininaya (myininaya):

lol sure male you just copied me ;)

OpenStudy (anonymous):

cos(x)-1

OpenStudy (anonymous):

LOL kk, sorry guys

OpenStudy (anonymous):

I think I am going to put "difference of squares" as a text shortcut on my computer so I can just paste it with a keypress...

OpenStudy (anonymous):

hello estudier. did you finish expected waiting time question?

OpenStudy (anonymous):

Haha estudier, too true. Around here at least xDD

OpenStudy (anonymous):

LOL - i know the diff of squares i'm just confused about the exponentiation convention with trig functions

OpenStudy (anonymous):

I think so, I am just trying to convince myself...

OpenStudy (anonymous):

@canuckish it has nothing at all to do with trig

myininaya (myininaya):

\[(trig(x))^2=trig^2(x)\]

OpenStudy (anonymous):

\[\frac{z^2-1}{z+1}=z-1\]

OpenStudy (anonymous):

@estudier i think problem is hard. maybe have to break up into even and odd tosses

OpenStudy (anonymous):

Oh yeah, same with the logarithms. \[(\cos(x))^n=\cos^n(x); (\ln(x))^n=\ln^n(x)\] Just don't confuse it with the arctrig functions. \[\frac{1}{\sin(x)}=\sin^{-1}(x)=\csc(x) \ne \arcsin(x)=\sin^{-1}(x)\] So you see the difference.

OpenStudy (anonymous):

i will go to post and tell you my thoughts

myininaya (myininaya):

honestly i never seen \[\ln^n(x)\]

OpenStudy (anonymous):

that's all my book does is square the function :( it's very confusing

myininaya (myininaya):

but i would think it would be equal to \[(\ln(x))^n\] if such a notation does exist

OpenStudy (anonymous):

Really? I use it all the time: \[\ln^2(x)\]

myininaya (myininaya):

\[\frac{1}{sinx}\neq \sin^{-1}x\]

OpenStudy (anonymous):

Well, kinda, just don't misinterpret it: \[\frac{1}{\sin(x)}=(\sin(x))^{-1}\]

OpenStudy (anonymous):

how are you guys getting the divisor symbol to show in the eq. builder? I just see (ab) and it doesn't have the line....

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ sinx\end{matrix}\right) = cscx ? yeah?\]

OpenStudy (anonymous):

see - no line :(

OpenStudy (anonymous):

\frac{a}{b}

OpenStudy (anonymous):

\[\frac{1}{sinx} = cscx\] there we go

OpenStudy (anonymous):

And yes it does equal that. :P